Compute g at a distance of 6. 8 x 10⁷m from the center of a spherical object whose mass is 5.0 x 10²³kg.


1
Expert's answer
2021-04-19T17:21:48-0400

Gravitational acceleration, created by the object of mass "M" at the distance "r" from the center of it is "g = G \frac{M}{r^2}", where "G = 6.67 \cdot 10^{-11} m^3 kg^{-1} s^{-2}" is gravitational constant.

Hence, "g = \frac{ 6.67 \cdot 10^{-11} m^3 kg^{-1} s^{-2} \cdot 5 \cdot 10^{23} kg }{(6.8 \cdot 10^7 m)^2} \approx 0.0072 \frac{m}{s^2}".